OpenStudy (anonymous):

how many real-number solutions does the equation have? –7x² + 6x + 3 = 0 one solution two solutions no solutions infinitely many solutions –8x² – 8x – 2 = 0 one solution two solutions no solutions infinitely many solutions

OpenStudy (mathstudent55):

Are you familiar with the quadratic formula?

OpenStudy (anonymous):

Kind of, I just don't really understand it.

OpenStudy (anonymous):

type khan academy quadratic equation into google. your questions will be answered.

OpenStudy (mathstudent55):

Here's the general form of a quadratic equation: ax^2 + bx + c = 0 If you compare your first quadratic equation, –7x² + 6x + 3 = 0, you'll notice that it follows the same form. In your case, a = -7, b = 6, and c = 3.

OpenStudy (mathstudent55):

The quadratic formula is a method of solving any quadratic equation of the form ax^2 + bx + c = 0 All you do is you substitute the values of a, b, and c into the formula to get the solutions to the quadratic equation.

OpenStudy (mathstudent55):